Solve for n.
(5 2/3 + 8 1/4) + 9 3/4 = n + 18

Answer by fractalier(6550)

From
(5 2/3 + 8 1/4) + 9 3/4 = n + 18
we can add 8 1/4 + 9 3/4 to get 18, which makes the problem simple
5 2/3 + 18 = n + 18
n = 5 2/3